I am having trouble with this word problem. Pleas, can someone help me?
Length of a rectangle is 2 inches more than twice its with. If the perimeter of the rectangle is 34 inches, find the dimension of the rectangle.
Thank you.

Answer by ptaylor(2198)

Let w=width of rectangle
Then 2w+2=length of rectangle
Perimeter=2w+2l=34 in
So we have
34=2w+2(2w+2) get rid of parens
34=2w+4w+4 subtract 4 from both sides
34-4=2w+4w+4-4 collect like terms
30=6w divide both sides by 6
w=5 in----------------------------width
2w+2=2*5+2=12 in-------------------length
CK
34=2*5+2*12
34=10+24
34=34

Hope this helps----ptaylor